OpenStudy (mendicant_bias):

Trying to solve another First order linear ODE using an integrating factor; workings posted below shortly.

OpenStudy (mendicant_bias):

\[2(1-x)y'-y=4x\sqrt{1-x}, \ \ \ y(0)=-1.\]

OpenStudy (mendicant_bias):

\[\frac{dy}{dx}-\frac{y}{2(1-x)}=\frac{4x \sqrt{1-x}}{2(1-x)}\]

OpenStudy (mendicant_bias):

\[\mu(x) \frac{dy}{dx}-\mu(x)\frac{y}{2(1-x)}=\mu(x) \frac{4x \sqrt{1-x}}{2(1-x)}\]

OpenStudy (mendicant_bias):

\[\mu(x)=e^{\int\limits_{}^{}P(x)dx}=e^{- \int\limits_{}^{}\frac{1}{2(1-x)}dx}\]

OpenStudy (mendicant_bias):

\[\frac{1}{2}\int\limits_{}^{}\frac{-1}{(1-x)}dx=\frac{1}{2}\ln(1-x) +C\] Getting rid of the constant of integration in this context,

OpenStudy (mendicant_bias):

\[\mu(x)=e^{\frac{1}{2}\ln(1-x)}=\sqrt{1-x}\]

OpenStudy (mendicant_bias):

\[f(x)\mu(x)=\frac{2x}{\sqrt{1-x}}\sqrt{1-x}=2x\]

OpenStudy (mendicant_bias):

\[y=\frac{\int\limits_{}^{}f(x)\mu(x)}{\mu(x)}dx=\frac{\int\limits_{}^{}2x \ dx}{\sqrt{1-x}}=\frac{x^2+C}{\sqrt{1-x}}\]

OpenStudy (mendicant_bias):

Sticking in the conditions applied to the IVP, \[y(0)=-1=\frac{(0)^2+C}{\sqrt{1-(0)}}=C; \ \ \ C = -1.\]

OpenStudy (mendicant_bias):

Answer is supposed to be:

ganeshie8 (ganeshie8):

Hint : \[x^2 - 1 = -(1-x^2) = -(1+x)(1-x)\]

OpenStudy (anonymous):

What they did was:\[ \frac{1-x}{\sqrt{1-x}} = \sqrt{1-x} \]

OpenStudy (mendicant_bias):

Yeah, I got it now. Thanks, nonetheless.